OpenStudy (sup_open~study):

What is the length of the hypotenuse of a right triangle with sides that measure 7 cm and 24 cm? A. 15 cm B. 25 cm C. 28 cm D. 35 cm

Parth (parthkohli):

Using the pythagorean theorem: \(\Large \color{MidnightBlue}{\Rightarrow a^2 + b^2 = c^2 }\) Where a,b are the legs and c is the hypotenuse. \(\Large \color{MidnightBlue}{\Rightarrow 7^2 + 24^2 = c^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 49 + 576 = c^2 }\) I'm gonna help you only till here.

Parth (parthkohli):

Hint: Add them and then find the square root.

Parth (parthkohli):

Considering you know how to solve equations. =)

OpenStudy (sup_open~study):

o now i remember thanks

OpenStudy (sup_open~study):

yes i do thank

Parth (parthkohli):

No problem dude ^-^